# GMLM: Bridging Graph Neural Networks and Language Models for Heterophilic Node Classification

### v3

@@ 4.3 Pre-trained Language Model (PLM) Branch with Active Node Selection
The PLM branch processes the textual descriptions associated with nodes to capture rich semantic information. To manage computational costs, particularly for large graphs, we incorporate an Active Node Selection mechanism. During each training iteration (in the fine-tuning stage), a subset of nodes $\mathcal{V}_{active}\subseteq\mathcal{V}$ is dynamically selected based on an active node mask $\mathbf{m}_{active}\in\{0,1\}^{N}$ (generation details in Section 5.2). Only the texts $\{t_{i}:v_{i}\in\mathcal{V}_{active}\}$ are processed by the PLM.
For each selected node $v_{i}\in\mathcal{V}_{active}$, its text $t_{i}$ is tokenized and fed into a pre-trained language models  [^1] We choose these models as they are computationally efficient but also maintain almost equivalent performance to their larger variants.: GTE-small[22] and DistilBERT [16]. The PLM outputs contextualized embeddings, from which we extract a fixed-size representation, specifically the embedding of the [CLS] token:
$\mathbf{h}_{PLM,i}=\text{PLM-Encoder}(t_{i})_{[CLS]}\in\mathbb{R}^{d_{PLM}}\quad\forall v_{i}\in\mathcal{V}_{active}$
For nodes $v_{j}\notin\mathcal{V}_{active}$, their PLM embeddings are effectively zero vectors when constructing the full PLM embedding matrix $\mathbf{H}_{PLM}^{full}\in\mathbb{R}^{N\times d_{PLM}}$ for fusion.
4.4 Fusion Network and Classification Head
To integrate the structural information from the GNN branch and the semantic information from the PLM branch, we employ a fusion network. For each node $v_{i}$, its projected GNN embedding $\mathbf{h}_{G,i}^{proj}$ (Equation 5) is concatenated with its corresponding (potentially zero) PLM embedding $\mathbf{h}_{PLM,i}^{full}$ from $\mathbf{H}_{PLM}^{full}$:
$\mathbf{h}_{concat,i}=[\mathbf{h}_{G,i}^{proj};\mathbf{h}_{PLM,i}^{full}]\in\mathbb{R}^{d_{PLM}+d_{PLM}}$ (7)
This concatenated vector is then passed through a fusion network, which is a multi-layer perceptron (MLP) designed to learn a joint representation:
$\mathbf{h}_{fused,i}=\text{MLP}_{fusion}(\mathbf{h}_{concat,i})\in\mathbb{R}^{d_{fused}}$ (8)
In our model, $\text{MLP}_{fusion}$ consists of two linear layers interspersed with Layer Normalization, ReLU activation, and Dropout. It projects the $2\cdot d_{PLM}$ dimensional concatenated vector to an intermediate hidden dimension, and then to a $d_{fused}$ dimensional fused representation, where $d_{fused}=d_{PLM}$. The final output of the fusion network also undergoes Layer Normalization.
@@ 4.4 Fusion Network and Classification Head
Finally, the fused representation $\mathbf{h}_{fused,i}$ is fed into a classification head, another MLP (Linear -> ReLU -> Dropout -> Linear), to predict the node label probabilities:
@@
where $\hat{\mathbf{y}}_{i}\in\mathbb{R}^{|\mathcal{Y}|}$ contains the predicted probabilities for each class in $\mathcal{Y}$.
5 GMLM Training Procedure
The training of GMLM proceeds in two distinct stages: (1) GNN Contrastive Pretraining, focused on learning robust structural representations and initializing the GNN-specific mask token, followed by (2) End-to-End Fine-tuning, where the entire model is trained for the downstream node classification task. All training stages and experiments for all 3 datasets were conducted on a single Tesla V100 (32GB) and took about $\sim$30 minutes with a maximum VRAM usage of 3GB. We also provide the impact on embeddings across different stages of training in  Appendix D.
5.1 Stage 1: GNN Contrastive Pretraining
The primary goal of this stage is to pre-train the GNN branch (Section 4.2) and the learnable GNN-specific mask token embedding $\mathbf{e}_{mask}^{gnn}$ (Section 4.1) to capture meaningful structural information from the graph. This is achieved through a contrastive learning objective. During this stage the weights of the PLM are frozen, but become trainable during Stage 2 as described in Section 5.2.
@@ 5.1.1 Data Augmentation via Soft Masking
For each training iteration in this stage, we generate two augmented "views" of the input node features. Two distinct perturbation masks, $\mathbf{m}_{pert,1}$ and $\mathbf{m}_{pert,2}$, are created by randomly selecting a proportion $p_{gnn\_mask}$ (randomly sampled from a uniform distribution $\mathcal{U}(0.2,0.5)$ in each iteration) of all nodes in the graph $\mathcal{V}$ to be perturbed. This selection can be uniform random or based on heuristics like node degree (the code uses degree-based sampling if degrees are non-zero, otherwise random). Using these two masks and the soft masking mechanism (Equation 4) with an interpolation weight $\beta$ ($\beta=0.7$), we obtain two sets of perturbed node features, $\mathbf{X}^{\prime}_{1}$ and $\mathbf{X}^{\prime}_{2}$.
@@ 5.1.2 Contrastive Objective
Both $\mathbf{X}^{\prime}_{1}$ and $\mathbf{X}^{\prime}_{2}$ are fed through the same shared-weight GNN encoder (specifically, we produce two sets of node embeddings, $\mathbf{H}_{G,1}$ and $\mathbf{H}_{G,2}\in\mathbb{R}^{N\times d_{G}}$. We then apply the NT-Xent contrastive loss (Equation 3) with a temperature $\tau_{pretrain}$ ($\tau_{pretrain}=0.5$). The pretraining loss $\mathcal{L}_{pretrain}$ is:
@@
We conduct contrastive pre-training for 30 epochs. During this pretraining phase, only the parameters of the GNN branch and $\mathbf{e}_{mask}^{gnn}$ are updated using an AdamW optimizer. Specifically, the GNN branch parameters (RGCN layers, batch normalization, residual projections) use a learning rate $lr_{gnn\_branch\_pretrain}$ ($1\times 10^{-3}$), while the GNN-specific mask token $\mathbf{e}_{mask}^{gnn}$ uses a learning rate $lr_{token\_pretrain}$ ($1\times 10^{-4}$). Both use a weight decay of $0.01$. A Cosine Annealing learning rate scheduler is also employed.
@@ 5.2 Stage 2: End-to-End Fine-tuning for Node Classification
After GNN pretraining, the entire GMLM architecture, with the GNN branch initialized from Stage 1, is fine-tuned for node classification. We utilize the raw textual information associated with heterophilic graphs as input to the PLM. Such information is only available for three datasets (Cornell, Texas, and Wisconsin) which we obtain from Zenodo  [^2] https://zenodo.org/records/7555368.
5.2.1 Active Node Selection for PLM and Loss Computation
For each training iteration, an active node mask $\mathbf{m}_{active}$ is generated. This mask first considers nodes in the current training split ($\mathcal{V}_{train}$). From this set, a random proportion $p_{active\_nodes}$ (randomly sampled from a uniform distribution $\mathcal{U}(0.3,0.8)$ in each iteration) is selected based on node degree (or randomly if all degrees are zero). Only texts $\{t_{i}:m_{active,i}=1\}$ are processed by the PLM, and the classification loss is computed only over these active nodes.
5.2.2 GNN Input Perturbation during Fine-tuning
The soft masking mechanism (Equation 4) with the same interpolation weight $\beta$ ($\beta=0.7$) is also applied to the GNN input features during fine-tuning. The perturbation mask used is $\mathbf{m}_{active}$, meaning nodes whose texts are processed by the PLM also have their GNN input features (partially) masked by $\mathbf{e}_{mask}^{gnn}$ (which is also being fine-tuned).
@@ 5.2.3 Supervised Classification Loss
The model processes inputs to produce predicted label probabilities $\hat{\mathbf{y}}_{i}$. The supervised classification loss $\mathcal{L}_{classify}$ is the cross-entropy loss computed over active nodes $\mathcal{V}_{active}$. Label smoothing (with a factor of 0.1) is applied to the true labels $y_{i,c}$.
@@ 5.2.4 Optimization and Hyperparameters
The total loss $\mathcal{L}_{classify}$ is backpropagated. All trainable parameters are updated using the AdamW optimizer [23]. We employ differential learning rates:
  • GNN components (RGCN layers, residual projections, batch normalization): $lr_{gnn}$ ($1\times 10^{-3}$).
  • PLM (GTE-small & DistilBERT) parameters: $lr_{plm}$ ($1\times 10^{-5}$).
@@
  • Other components (fusion network, classifier, GNN projection, GNN-specific mask token $\mathbf{e}_{mask}^{gnn}$): $lr_{other}$ ($1\times 10^{-4}$).
@@
A weight decay of $0.01$ is applied to all parameter groups. The model is trained for a maximum of 1000 epochs. We employ an early stopping strategy with a patience of 15 epochs: training halts if the F1 score on the validation set does not improve. The model state corresponding to the best validation F1 score is saved and used for final evaluation on the test set. The specific values for hyperparameters were set as follows (selected through a grid search the details of which are provided in  Appendix C):
  • $p_{gnn\_mask}$ (GNN Perturbation Ratio for Pretraining): During GNN contrastive pretraining, for each of the two generated views, the proportion of nodes whose input features are soft-masked is randomly sampled from a uniform distribution $\mathcal{U}(0.2,0.5)$ in each epoch.
@@
  • $p_{active\_nodes}$ (Active Node Ratio for Fine-tuning): In each epoch of the end-to-end fine-tuning stage, the proportion of training nodes selected to be "active" (for PLM processing, GNN input perturbation, and loss computation) is randomly sampled from a uniform distribution $\mathcal{U}(0.3,0.8)$.
  • $\beta$ (Soft Masking Interpolation Weight): The interpolation weight for the soft masking mechanism (Equation 4) is set to $\beta=0.7$. This value is used during both the GNN pretraining stage and for GNN input feature perturbation in the fine-tuning stage.
  • $\tau_{pretrain}$ (Pretraining Contrastive Temperature): The temperature hyperparameter for the NT-Xent contrastive loss (Equation 3) during the GNN pretraining stage is set to $\tau_{pretrain}=0.5$.
Cornell Wisconsin Texas
MLP-2 91.30 $\pm$ 0.70 93.87 $\pm$ 3.33 92.26 $\pm$ 0.71
GAT[4] 76.00 $\pm$ 1.01 71.01 $\pm$ 4.66 78.87 $\pm$ 0.86
APPNP[24] 91.80 $\pm$ 0.63 92.00 $\pm$ 3.59 91.18 $\pm$ 0.70
GPRGNN[25] 91.36 $\pm$ 0.70 93.75 $\pm$ 2.37 92.92 $\pm$ 0.61
H2GCN[26] 86.23 $\pm$ 4.71 87.5 $\pm$ 1.77 85.90 $\pm$ 3.53
MixHop[27] 60.33 $\pm$ 28.53 77.25 $\pm$ 7.80 76.39 $\pm$ 7.66
GCN+JK 66.56 $\pm$ 13.82 62.50 $\pm$ 15.75 80.66 $\pm$ 1.91
GAT+JK 74.43 $\pm$ 10.24 69.50 $\pm$ 3.12 75.41 $\pm$ 7.18
FAGCN [28] 88.03 $\pm$ 5.6 89.75 $\pm$ 6.37 88.85 $\pm$ 4.39
BernNet[29] 92.13 $\pm$ 1.64 NA 93.12 $\pm$ 0.65
GraphSAGE[30] 71.41 $\pm$ 1.24 64.85 $\pm$ 5.14 79.03 $\pm$ 1.20
Geom-GCN*[31] 60.81 64.12 67.57
SGC-1 [32] 70.98 $\pm$ 8.39 70.38 $\pm$ 2.85 83.28 $\pm$ 5.43
SGC-2 [32] 72.62 $\pm$ 9.92 74.75 $\pm$ 2.89 81.31 $\pm$ 3.3
GCNII [33] 89.18 $\pm$ 3.96 83.25 $\pm$ 2.69 82.46 $\pm$ 4.58
GCNII+ [33] 90.49 $\pm$ 4.45 89.12 $\pm$ 3.06 88.52 $\pm$ 3.02
GCN [34] 82.46 $\pm$ 3.11 75.5 $\pm$ 2.92 83.11 $\pm$ 3.2
Snowball-2 82.62 $\pm$ 2.34 74.88 $\pm$ 3.42 83.11 $\pm$ 3.2
Snowball-3 82.95 $\pm$ 2.1 69.5 $\pm$ 5.01 83.11 $\pm$ 3.2
GMLM(GTE-Small) 96.74 $\pm$ 0.72 91.54 $\pm$ 2.23 94.83 $\pm$ 1.31
GMLM(DistilBERT) 96.87 $\pm$ 0.57 92.32 $\pm$ 2.21 95.16 $\pm$ 1.27
Table 1: Mean Test Accuracy(%) $\pm$ standard deviation on three heterophilic node classification tasks across 10 runs. Highest score is denoted by bold and second highest by underline. $\ast$ Denotes the best performing variant of multiple variants reported by the authors.

### v4

@@ 4.3 Pre-trained Language Model (PLM) Branch
The PLM branch processes the textual descriptions associated with nodes to capture rich semantic information. To manage computational costs, especially on large graphs, we process texts in micro-batches. During each training iteration, an active node mask $\mathbf{m}_{active}\in\{0,1\}^{N}$ identifies a subset of nodes $\mathcal{V}_{active}$ whose texts will be processed.
For each node $v_{i}\in\mathcal{V}_{active}$, its text $t_{i}$ is tokenized and fed into a pre-trained language model, such as GTE-Base (110M)Li et al. (2023b) and Snowflake-Embed(305M)Yu et al. (2024). To derive a fixed-size representation from the PLM’s variable-length output, we perform mean pooling over the last hidden state, weighted by the attention mask to exclude padding tokens. The PLM embedding for an active node $v_{i}$ is:
$\mathbf{h}_{PLM,i}=\frac{\sum_{j=1}^{L}\mathbf{o}_{j}\cdot a_{j}}{\sum_{j=1}^{L}a_{j}}$ (8)
where $\mathbf{O}=\text{PLM-Encoder}(t_{i})$ is the last hidden state sequence of length $L$, $\mathbf{o}_{j}$ is the embedding of the $j$-th token, and $a_{j}$ is the corresponding value from the attention mask (1 for real tokens, 0 for padding). For inactive nodes $v_{k}\notin\mathcal{V}_{active}$, their PLM embeddings are set to zero vectors, resulting in a complete PLM embedding matrix $\mathbf{H}_{PLM}\in\mathbb{R}^{N\times d_{PLM}}$.
4.4 Bi-Directional Attention Fusion and Classification
To achieve a deep and meaningful integration of graph and text modalities, we move beyond simple concatenation and employ a bi-directional cross-attention mechanism. This allows the graph and text representations to mutually inform and refine one another. The full node embedding matrices, $\mathbf{H}_{G}$ (from Section 4.2) and $\mathbf{H}_{PLM}$ (from Section 4.3), serve as inputs.
The cross-attention mechanism computes two sets of attended features:
  1. 1.
    Graph-to-Text Attention: The PLM embeddings serve as queries to attend to the GNN embeddings, producing text-aware graph features, $\mathbf{H}_{G\to T}$.
  2. 2.
    Text-to-Graph Attention: The GNN embeddings serve as queries to attend to the PLM embeddings, producing structure-aware text features, $\mathbf{H}_{T\to G}$.
The standard scaled dot-product attention is used, where for a query matrix $\mathbf{Q}$, key matrix $\mathbf{K}$, and value matrix $\mathbf{V}$:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}$ (9)
The two attended representations are then concatenated for each node:
$\mathbf{h}_{concat,i}=[\mathbf{h}_{G\to T,i}\,;\,\mathbf{h}_{T\to G,i}]\in\mathbb{R}^{2\cdot d_{PLM}}$ (10)
This concatenated vector is passed through a fusion network, which is a feed-forward network consisting of a linear layer, LayerNorm, a GELU activation, and Dropout, to produce the final integrated representation:
$\mathbf{h}_{fused,i}=\text{FusionNet}(\mathbf{h}_{concat,i})\in\mathbb{R}^{d_{fused}}$ (11)
@@ 4.4 Bi-Directional Attention Fusion and Classification
Finally, this fused representation $\mathbf{h}_{fused,i}$ is fed into a classification head, another MLP (Linear -> GELU -> Dropout -> Linear), to predict the node label probabilities:
@@
where $\hat{\mathbf{y}}_{i}\in\mathbb{R}^{|\mathcal{Y}|}$ contains the predicted probabilities for each class in the label set $\mathcal{Y}$.
5 Training Procedure
The model training proceeds in two distinct stages: (1) GNN Contrastive Pre-training, focused on learning robust structural representations and initializing the GNN-specific mask token, followed by (2) End-to-End Fine-tuning, where the entire model is trained for the downstream node classification task. This two-stage approach allows the GNN to develop a strong structural foundation before being integrated with the powerful but computationally intensive language model. We also provide an analysis of the impact on embeddings across different stages of training in LABEL:app:imp-train.
5.1 Stage 1: GNN Contrastive Pre-training
The primary goal of this stage is to pre-train the multi-scale GNN branch (Section 4.2) and the learnable GNN-specific mask token embedding $\mathbf{e}_{mask}^{gnn}$ (Section 4.1) to capture meaningful structural information. This is achieved through a contrastive learning objective. During this stage, the weights of the PLM encoder, fusion network, and final classifier are frozen.
@@ 5.1.1 Data Augmentation via Soft Masking
For each training iteration, we generate two augmented "views" of the graph. Two distinct perturbation masks, $\mathbf{m}_{pert,1}$ and $\mathbf{m}_{pert,2}$, are created by selecting a proportion $p_{gnn\_mask}$ (randomly sampled from $\mathcal{U}(0.2,0.4)$ each iteration) of all nodes in the graph. This selection is weighted by node degree to prioritize more connected nodes. Using these masks and the soft masking mechanism (Equation 4) with an interpolation weight $\beta=0.7$, we obtain two sets of perturbed node features, $\mathbf{X}^{\prime}_{1}$ and $\mathbf{X}^{\prime}_{2}$.
@@ 5.1.2 Contrastive Objective
Both $\mathbf{X}^{\prime}_{1}$ and $\mathbf{X}^{\prime}_{2}$ are passed through the same shared-weight GNN encoder to produce two sets of multi-scale node embeddings, $\mathbf{H}_{G,1}$ and $\mathbf{H}_{G,2}$. We then apply the NT-Xent contrastive loss with a temperature $\tau_{pretrain}=0.1$. The pre-training loss is:
@@
This pre-training is conducted for 30 epochs. Only the parameters of the GNN branch (RGCN layers, graph normalization, residual projections, multi-scale fusion) and the mask token $\mathbf{e}_{mask}^{gnn}$ are updated. We use an AdamW optimizer with a learning rate of $1\times 10^{-4}$ for all trainable parameters in this stage and a weight decay of $0.05$. A CosineAnnealingWarmRestarts learning rate scheduler is employed to adjust the learning rate throughout pre-training.
@@ 5.2 Stage 2: End-to-End Fine-tuning for Node Classification
After GNN pre-training, the entire model, with the GNN branch initialized from Stage 1, is fine-tuned for node classification. The model is designed to leverage raw textual information associated with nodes, which is particularly valuable for navigating the complex relationships in heterophilic graphs.
5.2.1 Active Node Selection and GNN Perturbation
For each training iteration, an active node mask $\mathbf{m}_{active}$ is generated. This mask first considers nodes in the training split ($\mathcal{V}_{train}$). From this set, a random proportion $p_{active\_nodes}$ (sampled from $\mathcal{U}(0.3,0.8)$) is selected based on node degree. This active mask serves two purposes:
  1. 1.
    PLM Processing: Only the texts $\{t_{i}:m_{active,i}=1\}$ are processed by the PLM to manage computational load.
  2. 2.
    GNN Input Perturbation: The same mask $\mathbf{m}_{active}$ is used to apply soft masking (Equation 4) to the GNN’s input features, continuing to regularize the GNN during fine-tuning.
The final classification loss is also computed exclusively over these active nodes.
@@ 5.2.2 Supervised Classification Loss
The model processes the inputs to produce predicted label probabilities $\hat{\mathbf{y}}_{i}$ (Equation 12). The supervised classification loss $\mathcal{L}_{classify}$ is the cross-entropy loss with label smoothing (smoothing factor of 0.2) applied to the true labels.
@@ 5.2.3 Optimization and Hyperparameters
The total loss $\mathcal{L}_{classify}$ is backpropagated through the entire model. All trainable parameters are updated using the AdamW optimizer Loshchilov and Hutter (2019) with differential learning rates:
  • GNN components (RGCN layers, projections, etc.): $lr_{graph}=1\times 10^{-4}$.
  • PLM (GTE-Base) parameters: $lr_{bert}=1\times 10^{-5}$.
@@
  • Other components (attention, fusion, classifier, mask token): $lr_{other}=1\times 10^{-4}$.
@@
A weight decay of $0.05$ is applied to the GNN and "other" parameter groups, while the PLM uses a standard weight decay of $0.01$. The model is trained for a maximum of 500 epochs. A linear learning rate scheduler with a warm-up period corresponding to the first 10% of training steps is used. Additionally, gradient clipping with a maximum norm of 1.0 is applied to prevent exploding gradients. We employ an early stopping strategy with a patience of 30 epochs: training halts if the F1 score on the validation set does not improve. The model state with the best validation F1 score is used for the final evaluation.
The key hyperparameters are summarized below:
  • $p_{gnn\_mask}$ (Pre-training Perturbation Ratio): Sampled from $\mathcal{U}(0.2,0.4)$ in each pre-training epoch.
@@
  • $p_{active\_nodes}$ (Fine-tuning Active Node Ratio): Sampled from $\mathcal{U}(0.3,0.8)$ in each fine-tuning epoch.
  • $\beta$ (Soft Masking Weight): Set to $0.7$ for both stages.
  • $\tau_{pretrain}$ (Pre-training Contrastive Temperature): Set to $0.1$.
Methods Cornell Texas Wisconsin Actor Amazon
Classic GNNs
GCN Kipf and Welling (2017) 52.86$\pm$1.8 43.64$\pm$3.3 41.40$\pm$1.8 66.70$\pm$1.3 39.33$\pm$1.0
GraphSAGEHamilton et al. (2017) 75.71$\pm$1.8 81.82$\pm$2.5 80.35$\pm$1.3 70.37$\pm$0.1 46.63$\pm$0.1
GAT 54.28$\pm$5.1 51.36$\pm$2.3 50.53$\pm$1.7 63.74$\pm$6.7 35.12$\pm$6.4
Heterophily-specific GNNs
H2GCNZhu et al. (2020b) 69.76$\pm$3.0 79.09$\pm$3.5 80.18$\pm$1.9 70.73$\pm$0.9 47.09$\pm$0.3
FAGCNBo et al. (2021) 76.43$\pm$3.1 84.55$\pm$4.8 83.16$\pm$1.4 75.58$\pm$0.5 49.83$\pm$0.6
JacobiConvWang and Zhang (2022) 73.57$\pm$4.3 81.80$\pm$4.1 76.31$\pm$11.3 73.81$\pm$0.3 49.43$\pm$0.5
GBK-GNNDu et al. (2022) 66.19$\pm$2.8 80.00$\pm$3.0 72.98$\pm$3.3 72.49$\pm$1.0 44.90$\pm$0.3
OGNNWang et al. (2022) 71.91$\pm$1.8 85.00$\pm$2.3 79.30$\pm$2.1 72.08$\pm$2.4 47.79$\pm$1.6
SEGSL 66.67$\pm$4.1 85.00$\pm$2.0 79.30$\pm$1.8 72.73$\pm$0.8 47.38$\pm$0.2
DisamGCLZhao et al. (2024) 50.48$\pm$2.0 65.00$\pm$1.2 57.89$\pm$0.0 67.78$\pm$0.3 43.90$\pm$0.4
LLM4HeG Wu et al. (2025) (fine-tuned LLM/SLMs and distilled SLMs)
Vicuna 7B 77.62$\pm$2.9 89.09$\pm$3.3 86.14$\pm$2.1 76.82$\pm$0.5 51.53$\pm$0.4
Bloom 560M 75.48$\pm$2.1 80.00$\pm$4.0 86.49$\pm$1.9 76.16$\pm$0.6 51.52$\pm$0.5
Bloom 1B 75.71$\pm$1.4 83.86$\pm$2.8 83.86$\pm$1.7 74.99$\pm$0.5 52.33$\pm$0.6
7B-to-560M 75.00$\pm$4.0 88.18$\pm$2.2 87.19$\pm$2.5 75.78$\pm$0.2 51.51$\pm$0.4
7B-to-1B 77.38$\pm$2.7 88.18$\pm$4.0 86.14$\pm$1.5 75.37$\pm$0.9 51.58$\pm$0.4
GMLM
GTE-Base 74.25$\pm$2.5 85.90$\pm$4.9 89.81$\pm$2.5 73.89$\pm$2.1 50.64$\pm$0.7
Snowflake-embed 79.16$\pm$4.8 97.18$\pm$2.9 92.04$\pm$1.6 76.81$\pm$0.9 54.63$\pm$0.4
Table 1: Accuracy for node classification of different methods. Bold denotes the best score while underline denotes the second best.